OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). -1 1 0 1/2

OpenStudy (anonymous):

gradient is (y2-y1)/(x2-x1) = (-5--3)/(3-1) = -2/2 - -1

jhonyy9 (jhonyy9):

so when you have there (1,-3) what can you deduct from this ? - how many is x and y ?